Problem 12.16 Question Help Abey Kuruvilla, of Parkside Plumbing, uses 1,200 of a certain spare part that costs $27 for eachB) what is the economic order quantity? units
Answer #1

Annual Demand = D = 1200 units

Ordering cost = S = $27

Annual Holding cost = H = $24

a) Order size Q1 = 25

Total Cost = \frac{D}{Q1}*S + \frac{Q1}{2}*H = \frac{1200}{25}*27 + \frac{25}{2}*24 = 1296 + 50 = $1346

Order size Q2 = 40

Total Cost = \frac{D}{Q2}*S + \frac{Q2}{2}*H = \frac{1200}{40}*27 + \frac{40}{2}*24 = 810 + 480 = $1290

Order size Q3 = 50

Total Cost = \frac{D}{Q3}*S + \frac{Q3}{2}*H = \frac{1200}{50}*27 + \frac{50}{2}*24 = 648 + 600 = $1248

Order size Q4 = 60

Total Cost = \frac{D}{Q4}*S + \frac{Q4}{2}*H = \frac{1200}{60}*27 + \frac{60}{2}*24 = 540 + 720 = $1260

Order size Q5 = 100

Total Cost = \frac{D}{Q5}*S + \frac{Q5}{2}*H = \frac{1200}{100}*27 + \frac{100}{2}*24 = 324 + 1200 = $1524

b) Economic Order quantity = \sqrt{\frac{2*D*S}{H}} = \sqrt{\frac{2*1200*27}{24}} = 51.96 units

-----------------------------------------------------------------------------------------------------------------------

In case of any doubt, please ask through the comment section before Upvote/downvote.

If you liked the answer, please give an upvote. This will be quite encouraging for me, thank-you!!!